# Calibrating Translation Decoding with Quality Estimation on LLMs

### v2

@@ Appendix A Detailed Dataset Description
Appendix B More Results for Section 5.1
In this appendix, we provide additional results from the experiments presented in the main text, including (1) a broader range of translation evaluation metrics, (2) the experimental results based on the ALMA model, and (3) the experimental results under on-policy training dynamics.
B.1 Off-policy Results based on Tower in KIWI-XXL and COMET
@@
B.2 Off-policy Results based on ALMA
Table 4: Out-of-English translation evaluation results for ALMA models on the WMT22 dataset, measured by 3 metrics, i.e., COMET, KIWI-XXL, and XCOMET. When applying our calibration method on ALMA-base, consistent substantial improvements across all metrics and directions can be observed.
ende encs COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 85.45 80.70 96.49 89.05 82.06 90.82 + SFT on preferred data 85.42 80.44 96.26 89.11 81.28 90.26 + DPO 85.19 80.02 96.22 88.78 81.03 90.12 + CPO (ALMA-13B-R) 86.06 82.77 97.11 89.61 84.81 91.91 + Calibration (ours) 86.22 83.83 97.25 90.24 86.25 93.06 enis enzh COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 85.44 81.51 89.94 84.87 77.14 88.11 + SFT on preferred data 85.19 80.25 89.15 85.36 78.16 88.34 + DPO 85.20 80.42 88.97 84.73 76.96 87.72 + CPO (ALMA-13B-R) 85.80 82.35 89.63 85.89 81.79 89.55 + Calibration (ours) 86.29 84.11 91.62 86.47 83.18 90.75 enru Avg. COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 87.05 82.60 92.98 86.37 80.80 91.67 + SFT on preferred data 86.88 81.79 92.57 86.39 80.38 91.32 + DPO 86.70 81.61 92.53 86.12 80.01 91.11 + CPO (ALMA-13B-R) 87.86 84.97 94.15 87.04 83.34 92.47 + Calibration (ours) 88.41 86.30 94.46 87.53 84.73 94.43
Table 5: Into-English translation evaluation results for ALMA models on the WMT22 dataset, measured by 3 metrics, i.e., COMET, KIWI-XXL, and XCOMET. When applying our calibration method on ALMA-base, consistent substantial improvements across all metrics and directions can be observed.
@@ B.2 Off-policy Results based on ALMA
Tables 4 and 5 present the experimental results based on the ALMA model under off-policy settings. Consistently substantial improvements across metrics — COMET-22 (abbreviated as COMET), KIWI-XXL, and XCOMET-XXL (abbreviated as XCOMET) — are observed when applying our calibration method. The resulting performance surpasses all baselines across nearly all language directions, including the strongest ALMA variant, ALMA-7B-R. For example, we observe gains of +0.5 COMET, +1.4 KIWI-XXL, and +2.0 XCOMET points on out-of-English translation compared to ALMA-7B-R.
B.3 On-policy Results based on Tower
@@
In this appendix, we provide additional results for Section C presented in the main text, including (1) another evaluation direction, namely, enru, and (2) another statistic, Kendall’s $\tau$ for metric meta-evaluation.
Figure 6 shows the Spearman coefficient and the corresponding translation performance in the enru direction. Meanwhile, Figures 7 and 8 present the results using Kendall’s $\tau$ for the ende and enru directions, respectively. It is clear that the main findings, as mentioned in Section 5.2, hold across language directions and statistics.
Appendix D Human Study
In this appendix, we present the detailed human evaluation results for three additional target languages: German, Russian, and Dutch. Section D.1 outlines the annotation criteria used, and Section D.3 reports the score distributions and aggregated summaries for these language directions.
D.1 Annotation Criteria
Note that for WMT, a continuous score is allowed for annotators. However, we require the annotator to score with integers. For example, a score of 5 means in between of minor error and perfection, which may be reflected in tone, grammar, or some subtle wording.
D.2 English-to-Chinese Translation Annotation Analysis
Note that annotators were blinded to the source of each translation to prevent any systematic bias.
D.3 Other Annotation Results
Annotation results in other translation directions are in preparation and will be published soon. Also, we will release all detailed human annotations to the community [^5] Detailed human annotations are available at https://huggingface.co/datasets/Calibration-Translation/Calibration-translation-human-eval..
Appendix E Prompt Templates
E.1 GPT Prompt
E.2 Tower Prompt
E.3 ALMA Prompt

### v3

@@ Appendix D Detailed Dataset Description
Appendix E Prompt Templates
E.1 GPT Prompt
E.2 Tower Prompt
E.3 ALMA Prompt
Appendix F More Results for Section 5.1
In this appendix, we provide additional results from the experiments presented in the main text, including (1) a broader range of translation evaluation metrics in Section F.1, (2) the experimental results based on the ALMA model in Section F.2, and (3) the experimental results under on-policy training dynamics in Section F.3.
F.1 Off-Policy Results based on Tower in KIWI-XXL and COMET
F.2 Off-Policy Results based on ALMA
Tables 4 and 5 present the experimental results based on the ALMA model under off-policy settings. Consistently substantial improvements across metrics—COMET-22 (abbreviated as COMET), CometKiwi-XXL (abbreviated as KIWI-XXL), and XCOMET-XXL (abbreviated as XCOMET)—are observed when applying our calibration method. The resulting performance surpasses all baselines across nearly all language directions, including the strongest ALMA variant, ALMA-7B-R. For example, we observe gains of +0.5 COMET, +1.4 KIWI-XXL, and +2.0 XCOMET points on out-of-English translation compared to ALMA-7B-R.
F.3 On-Policy Results based on Tower
Appendix G More Results for Section 5.2
In this appendix, we provide additional results complementing Section 5.2 of the main text, including (1) an additional evaluation direction, enru, and (2) an alternative statistic for metric meta-evaluation, Kendall’s $\tau$.
Figure 8 shows the Spearman coefficient and the corresponding translation performance in the enru direction. Meanwhile, Figures 9 and 10 present the results using Kendall’s $\tau$ for the ende and enru directions, respectively. It is clear that the main findings, as mentioned in Section 5.2, hold across language directions and statistics.
Appendix H Human Study
In this appendix, we present the detailed human evaluation results for three additional target languages: German, Russian, and Dutch. Section H.1 outlines the annotation criteria used, and Section H.3 reports the score distributions and aggregated summaries for these language directions.
H.1 Annotation Criteria
Note that for WMT, a continuous score is allowed for annotators. However, we require annotators to score with integers. For example, a score of 5 means in between of minor error and perfection, which may be reflected in tone, grammar, or some subtle wording.
H.2 English-to-Chinese Translation Annotation Analysis
Note that annotators were blinded to the source system of each translation by shuffling the order of each translation pair to prevent systematic (order) bias.
H.3 Other Annotation Results
Human Evaluation Summary
TScore CScore Win Loss Tie
4.79 5.21 93 44 63
Figure 7: Human evaluation results in enru direction: score distribution and an aggregated summary. TScore and CScore denote the average scores over 200 annotations for TowerInstruct-Mistral-7B and our calibrated model, respectively. Win, Loss, and Tie indicate the number of samples for which our model outperformed, underperformed, or tied with TowerInstruct-Mistral-7B.
We will release all detailed human annotation results in other translation directions to the community [^5] Detailed human annotations are available at https://huggingface.co/datasets/Calibration-Translation/Calibration-translation-human-eval.
@@ H.3 Other Annotation Results
Table 4: Out-of-English translation evaluation results for ALMA models on the WMT22 dataset, measured by COMET, KIWI-XXL, and XCOMET. When applying our calibration method on ALMA-base, consistent substantial improvements across all metrics and directions can be observed.
ende encs COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 85.45 80.70 96.49 89.05 82.06 90.82 + SFT on preferred data 85.42 80.44 96.26 89.11 81.28 90.26 + DPO 85.19 80.02 96.22 88.78 81.03 90.12 + CPO (ALMA-7B-R) 86.06 82.77 97.11 89.61 84.81 91.91 + Calibration (ours) 86.22 83.83 97.25 90.24 86.25 93.06 enis enzh COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 85.44 81.51 89.94 84.87 77.14 88.11 + SFT on preferred data 85.19 80.25 89.15 85.36 78.16 88.34 + DPO 85.20 80.42 88.97 84.73 76.96 87.72 + CPO (ALMA-7B-R) 85.80 82.35 89.63 85.89 81.79 89.55 + Calibration (ours) 86.29 84.11 91.62 86.47 83.18 90.75 enru Avg. COMET KIWI-XXL XCOMET COMET KIWI-XXL XCOMET ALMA-7B-LoRA 87.05 82.60 92.98 86.37 80.80 91.67 + SFT on preferred data 86.88 81.79 92.57 86.39 80.38 91.32 + DPO 86.70 81.61 92.53 86.12 80.01 91.11 + CPO (ALMA-7B-R) 87.86 84.97 94.15 87.04 83.34 92.47 + Calibration (ours) 88.41 86.30 94.46 87.53 84.73 94.43
Table 5: Into-English translation evaluation results for ALMA models on the WMT22 dataset, measured by COMET, KIWI-XXL, and XCOMET. When applying our calibration method on ALMA-base, consistent substantial improvements across all metrics and directions can be observed.